# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v3

@@ 2.1 Bayesian 3D modeling
Figure 4 reveals the key characteristics of the individual PDFs of $\Gamma$ for a set of evenly sampled data. First, in a majority of the 20 cases, the peaks and medians of individual PDFs tend to be biased toward values lower than the input value (i.e. $\Gamma=0$ in the present case) and the tails go to zero more slowly in the negative direction than in the positive direction, while the opposite is true in the rest of the minority cases. Second, each individual PDF tends to be broader in the majority cases than in the minority cases. Finally, due to the said properties of the individual PDFs the averaged PDF (the thick black curve) has a median biased toward a negative value. This bias is particularly pronounced for the data set of the $a=8$ kau binary.
The bias (i.e. the failure to reproduce the input value) of the averaged PDF indicates that it is a poor quantity to combine individual PDFs. The averaged PDF has another fatal property as a combination of independent results. Its width is broader than individual PDFs meaning that many independent observations/experiments did not lead to any improvement in the measurement of $\Gamma$. This does not agree with common sense knowledge of the increased precision from $N$ independent measurements.
@@
whose uncertainty follows from the propagation of the uncertainties of $x_{i}$, i.e. $\sigma_{i}$, as
@@
For an ideal case of $x_{i}=\mu$ and $\sigma_{i}=\sigma$ (i.e. all measurements return the same value and the same uncertainty), we have $\mu^{\prime}=\mu$ and $\sigma_{\mu^{\prime}}=\sigma/\sqrt{N}$. Thus, as is well known, the rule of thumb is that with $N$ independent measurements of the same quantity the precision is improved by a factor of $\sqrt{N}$.
Given that the averaged PDF is not satisfactory, what would be the best way to extract the maximal information from the set of individual PDFs, or how to combine independent measurements of the same quantity? This is the question statisticians have asked and come up with reasonable ideas. One particular suggestion to follow here is the consolidation by conflation [^4] Throughout I will refer to this as “consolidation by conflation” or simply “consolidation”. due to T. P. Hill and J. Miller (Hill, 2011; Hill & Miller, 2011). In this approach the consolidated probability distribution of a variable $x$ is defined by
@@
Lacking general analytic results in general non-Gaussian situations and considering the unknown nature of the present problem, here an empirical approach is followed, i.e., the consolidated distribution (Equation (30)) is tried for mock individual PDFs and its property is experimentally investigated. Figure 4 shows examples of such a numerical experiment for the mock data sets of the two binaries shown in Figure 2. Remarkably, the consolidated distributions are close to Gaussian in both cases despite individual non-Gaussianities. As will be shown later, the (near) Gaussianity of the consolidated distribution is always satisfied empirically for any sufficiently large set of individual PDFs of mock or real binaries. Figure 4 indicates that the (near) Gaussianity is a consequence of the product (i.e. the conflation by Equation (30)) of the shapes of the individual PDFs (which were evenly sampled in $t/P$). As the right column of the figure shows the dramatic nature, although the majority of individual PDFs are biased toward the negative side (which is why the averaged PDF is also biased as discussed above), the minority rest is biased toward the positive side with narrower widths so as to make a “correct” balance with the majority.
In addition to being (nearly) Gaussian, the consolidated distribution correctly reproduces the input value of $\Gamma=0$. As indicated in Figure 4, the consolidated distributions give $\Gamma=0.005\pm 0.026$ and $-0.004\pm 0.042$ for the two sets of data. Note that the former value is more precise because the relative uncertainties of $v_{r}$ are smaller for the data of the binary with $a=8$ kau for the fixed uncertainty of $\sigma_{v_{r}}\approx 283$ m s-1. Although not shown here, precision can also be improved by increasing the sample size. These numerical experiments suggest that the consolidated distribution of $\Gamma$ from a sufficiently well sampled individual PDFs have the potential to discriminate between Newtonian gravity and nonstandard gravities.
@@
In the present work, only the gravity anomaly parameter $\Gamma$ is of interest and can be consolidated for real data of wide binaries because the same gravity law is assumed in a certain acceleration range, but snapshot-observed wide binaries have different masses, orbital parameters, and inclinations, all of which are treated as nuisance parameters. However, the two specific sets of mock data shown in Figure 2 are special because each data set is composed of data for the same binary observed at sufficiently different times $t/P$, which of course cannot be realized in actual observations of wide binaries for obvious reasons. [^5] Not to mention, such a data set can be obtained for very close binaries with $P$ sufficiently small, e.g. several years. It is of some interest to consider the consolidation of the nuisance parameters for these mock data sets, not only as a matter of additionally testing the consolidation method, but also as a possible application of the method in very close binaries.
@@
Similarly, Figure 6 shows that the individual PDFs of eccentricity $e$ are not well shaped in general, but the consolidated distributions are well localized, nearly Gaussian, and reproduce the input value $e=0.75$ remarkably well.
Figure 7 shows the results for $\phi_{0}$. Because $\phi_{0}$ was allowed to vary freely in the range $-360^{\circ}<\phi_{0}<360^{\circ}$ and the input value was $-135^{\circ}$, we expect that individual PDFs will be localized around $-135^{\circ}$ or $225^{\circ}$. Most PDFs are indeed so, but there are a few exceptions. The most pronounced are the PDFs for $t/P=0.90$ and $0.95$, which correspond to the two points on the second quadrant of the $x^{\prime}y^{\prime}$ plane in Figure 2. Due to these outliers, the consolidated distributions (after transforming PDFs localized around $225^{\circ}$ by $-360^{\circ}$) are biased by $\approx-16^{\circ}$ and $\approx-22^{\circ}$ respectively for the two data sets of $a=0.8$ and $a=8$ kau, as can be seen in the lower panels. Compared with the net allowed range of $360^{\circ}$ the magnitude of these biases in $\phi_{0}$ is not large.
The bias in the PDFs of $\phi_{0}$ can be understood as follows. Equations (10) and (12) have dependence on $\phi_{0}$, but when their observational constraints given by Equations (9) and (6) are not strong enough, the prior constraint given by Equation (19) will play a major role. This means that some data near the periastron (that is, when $\phi\sim\phi_{0,\rm{true}}$) can lead to PDFs localized near $\phi_{0}\sim\phi\pm 180^{\circ}\sim\phi_{0,\rm{true}}\pm 180^{\circ}$ leading to bias. This appears to be the case for the two above-mentioned outliers in each data set.
Thus, unlike $i$ and $e$, $\phi_{0}$ appears to be not perfectly recovered by the procedure for the above mock observations. A couple of relevant remarks need to be made here. First, as Figure 4 shows, the relatively small biases in $\phi_{0}$ do not affect the correct recovery of $\Gamma$ in both data sets. This can be understood by the fact that only Equation (10) has the direct dependence on $\Gamma$, and therefore the small bias in $\phi_{0}$ has little impact. Second, the bias is larger when the relative precision of the radial velocities is poorer, i.e., for the data set with $a=8$ kau. It is interesting to check whether the bias is reduced if the assumed uncertainties of the radial velocities are smaller for the same data. Figure 8 shows the PDFs of the four parameters for the mock binary with $a=8$ kau, but assuming 50 m s-1 for RV uncertainties so that $\sigma_{v_{r}}\approx\sqrt{2}\times 50\approx 71$ m s-1 for the relative radial velocity $v_{r}$. The bias in $\phi_{0}$ is now reduced to $\approx-14^{\circ}$, while $\Gamma$ and the other parameters remain recovered correctly with smaller uncertainties as expected.
@@
Therefore, numerical experiments indicate that the procedure of Bayesian inferences with their consolidation can be used to measure $\Gamma$. Moreover, for a data set of one binary as in close binaries, the procedure can recover the parameters $e$ and $i$ correctly, and $\phi_{0}$ as well if the relative precision of velocities is good enough, which can be satisfied for close binaries. In close binaries, gravity is Newtonian (i.e. $\Gamma=0$ is fixed), so the procedure can be used to measure the total mass of the binary.
The above numerical experiments are designed to illustrate individual PDFs of parameters and develop/test the method of consolidation. For this reason, the mock data used in the above experiments have accurate values of velocities predicted by the Newtonian model. The assumed uncertainties of the velocities are only used to derive the PDFs of the model parameters. However, measured velocities will have errors due to measurement uncertainties, and thus errors need to be added to the model-predicted velocities for realistic simulations. For wide binary samples derived from the Gaia DR3 database to be considered in the next subsection, the projected components $v_{x^{\prime}}$ and $v_{y^{\prime}}$ have negligibly small errors, but the radial velocity $v_{r}$ has large errors. Realistic simulations with the errors of $v_{r}$ added to the model predicted values will be considered when the results of Bayesian modeling for real data are presented in Section 3.
@@ 2.3 Binary Sample
This study requires a sample of pure binaries that are free of hidden close components. Selecting and using pure binaries was originally considered by Hernandez et al. (2022) and Hernandez (2023) in the context of statistical tests of gravity with $v_{p}$. Here I apply a multi-step screening process to ensure that no (or extremely few) kinematic contaminated systems survive. I start with the initial selection criteria that were used by Chae (2024a, b). The initial selection is based on the requirement that binaries in a bin of sufficiently high accelerations must agree with the expectation of Newtonian dynamics in a statistical sense, specifically in terms of the median or average properties. Such a selection (Chae, 2024a, b) can be achieved by requiring sufficiently good precision of proper motions (PMs), parallaxes, and radial velocities (RVs) for both components and dynamical consistency between them. Note that availability of Gaia measured RVs (with a reasonable precision) particularly helps removing systems with hidden close companions because it implies a good single-star solution for the spectroscopic observations. However, as already acknowledged in Chae (2024a, b), there can be (a small number of) individual exceptions that elude such a statistical selection. [^6] Minor bumps in the high-velocity tails of normalized projected velocity distributions in Figures 15 - 17 of Chae (2024c) may indicate such exceptions.
Here I include some of $\delta_{A}<-28^{\circ}$ (hereafter $\delta$ refers to the declination and the subscript $A$ refers to the brighter component of the pair) binaries as opposed to Chae (2023a, 2024a) who excluded the $\delta_{A}<-28^{\circ}$ data altogether (if distance is greater than 80 pc) because the dust extinction model adopted by Chae (2023a) did not cover that portion of the southern sky. I include those binaries with $\delta_{A}<-28^{\circ}$ that satisfy either distance $d_{A}<100$ pc or galactic latitude $|b|>60^{\circ}$ as they are unlikely to be significantly affected by dust extinction (see Figure 3 of Chae 2023a). Another minor change is that I now select stars with $13.43.8<M_{G}<$ 13.4 (rather than $144<M_{G}<$ 14 where $M_{G}$ is the Gaia $G$-band absolute magnitude) in agreement with Figure 1 of Penoyre et al. (2022b).
@@
Note. (1) The base sample is defined as a sample with $f_{\rm{multi}}\rightarrow 0$ ($f_{\rm{multi}}$ is the fraction of hierarchical systems) in the Newtonian regime (it may include individual exceptions of hierarchical systems). This sample is revised and extended from Chae (2024a, b) to include some southern-sky binaries. (2) The base sample was selected from the El-Badry et al. (2021) catalog. Thus, it is free from fly-by (chance alignment) and resolved triple (or higher-order) cases by their criteria. Here I use stricter criteria to remove any potentially suspicious systems as described in the text. Out of 4276 systems, there are 68 potential fly-bys and 35 triple+ systems. (3) The nominal sample is used to obtain nominal results, and the alternative sample and other samples with varied ruwe and $\sigma_{v_{r}}$ limits are considered to explore possible systematic uncertainties.
The above selection returns a sample of 4,276 wide binaries that is larger than the Chae (2024b) sample of 3,557 wide binaries thanks to the inclusion of some of the southern-sky binaries. This sample will be referred to as the ‘base sample’. Table 1 gives a summary of the definition of the samples considered in this work, including the base sample. This sample is free of fly-bys (chance alignments) and resolved triples or higher-order multiples (to be referred to as triple+) by the criteria of El-Badry et al. (2021). Here I consider a stronger criterion to remove any potentially suspicious systems because individual modeling requires individually pure binaries. The neighboring space around each binary is defined by the following limits of the transverse and longitudinal separations:
@@
    The sky-projected displacement from the brighter component satisfies $\Delta R_{\rm{sky}}<\sqrt{20}s$ where $s$ is the projected separation between the pair.
@@
I search the entire Gaia catalog of all objects within 200 pc downloaded from the Gaia DR3 data archive [^7] https://gea.esac.esa.int/archive/. For 35 systems out of the base sample, one or more stars are found beyond the known pair within the space. I identify these as potential resolved triple+ systems. For 68 systems, the fainter component of the pair as originally identified by El-Badry et al. (2021) is not found within the space defined above. I identify these as potential fly-bys. In total, 103 (2.4%) of the 4276 systems have potential kinematic contaminants. I flag these systems in the base sample and exclude them in the following sample selection.
@@
This work requires accurate and precise values for all three components of the relative velocity between the pair. This requirement is met for the two components of the projected velocity $v_{p}$ from the Gaia DR3 database: most projected velocities have uncertainties $\sigma_{v_{p}}\lesssim 35$ m s-1 as shown in Figure 9. Here $v_{p}$ and $\sigma_{v_{p}}$ are calculated using Equations (5) and (9) of Chae (2024a). Radial velocities $v_{r}$ from the Gaia DR3 database have much larger uncertainties $\sigma_{v_{r}}>168$ m s-1. The uncertainty of $v=|\mathbf{v}|=\sqrt{v_{p}^{2}+v_{r}^{2}}$ ($\sigma_{v}$) is estimated from Monte Carlo distributions of $v_{p}$ and $v_{r}$ with their estimated uncertainties ($\sigma_{v_{p}}$ and $\sigma_{v_{r}}$) assuming normal distributions.
As $\sigma_{v_{r}}$ increases, $\sigma_{v}$ also increases and individual Bayesian modeling becomes increasingly less feasible and model parameters become more poorly constrained (and thus the results become less meaningful). As shown in Figure 2, wide binaries of low internal acceleration have relative velocities between the pair as low as a few hundreds m s-1 or lower. Thus, for wide binaries of low internal acceleration, $\sigma_{v_{r}}$ significantly larger than a few hundreds m s-1 are nor likely to be useful. I put an initial hard cut $\sigma_{v_{r}}<500$ m s-1, which leaves only 652 binaries (15.6% of the base sample). This will be referred to as the ‘statistical sample’ for Bayesian 3D modeling of this work.
Because reasonable precision is required for $v_{r}$ and a significant number of binaries are required for an unbiased inference of gravity (as will be shown below in Section 3), some compromise between the sample size and the tolerance of $\sigma_{v_{r}}$ is inevitable. Based on numerical experiments, I take $380$ m s-1 as a nominal choice for the upper limit of $\sigma_{v_{r}}$ but also consider $330$ m s-1 as an alternative choice. For $\sigma_{v_{r}}$ limits higher than $380$ m s-1, the bias caused by RV scatters seems to outweigh the increase in number statistics. For $\sigma_{v_{r}}$ limits lower than $330$ m s-1, the number statistics in the MOND regime is poor.
@@
The statistical sample with relatively precise $v_{r}$ values may still contain exceptional cases with kinematic contamination due to, e.g., unseen low-mass star(s) or massive Jovian planet(s) orbiting any component(s) of the binary system. A popular indicator of such close kinematic contamination is the Gaia renormalized unit weight error (ruwe) parameter [^8] https://gea.esac.esa.int/archive/documentation/. For general stars in the Gaia DR3 database a value significantly greater than $1.0$ (e.g., ruwe $>1.4$) could indicate that the source is non-single or otherwise problematic for the astrometric solution. Figure 10 shows the distribution of ruwe values for 1304 stars in the statistical sample; 92.2% (82.6%) satisfy ruwe $<1.4$ ($<1.2$). Nevertheless, it is possible that some of stars with ruwe $<1.4$ (or some other value) are unresolved non-single stars.
@@
Various simulations and observational studies suggest that ruwe values ranging from $1.15$ $1.6$ (e.g. Belokurov et al. 2020; Bryson et al. 2020; Fitton et al. 2022; Penoyre et al. 2022a; Castro-Ginard et. al. 2024) can help remove some (unresolved) close binaries. Since such a ruwe value does not provide a hard cut on close binaries (see, e.g., Figure 1 of Fitton et al. 2022) and the statistical sample is purer than the base sample (which is already pure in a statistical sense), I consider varying the upper limit of allowed ruwe values to explore the possible effects of the ruwe limit. For this purpose, I take $1.30$ as the nominal/standard choice for the upper limit of ruwe (see Figure 10). I consider other limits from ruwe $=1.1$ $2.0$ to explore possible systematic effects. Based on the nominal choices of both $\sigma_{v_{r}}$ and ruwe, I define the nominal/standard sample of 312 binaries as given in Table 1.

### v4

@@ 2.1 Bayesian 3D modeling
Figure 4 reveals the key characteristics of the individual PDFs of $\Gamma$ for a set of evenly sampled data. First, in a majority of the 20 cases, the peaks and medians of individual PDFs tend to be biased toward values lower than the input value (i.e., $\Gamma=0$ in the present case) and the tails go to zero more slowly in the negative direction than in the positive direction, while the opposite is true in the rest of the minority cases. Second, each individual PDF tends to be broader in the majority of cases than in the minority of cases. Finally, due to the said properties of the individual PDFs the averaged PDF (the thick black curve) has a median biased toward a negative value. This bias is particularly pronounced for the data set of the $a=8$ kau binary.
The bias (i.e., the failure to reproduce the input value) of the averaged PDF indicates that it is a poor way of combining individual PDFs. The averaged PDF has another fatal property as a way of combining independent results. Its width is broader than individual PDFs meaning that many independent observations/experiments did not lead to any improvement in the measurement of $\Gamma$. This does not agree with common sense knowledge of the increased precision from $N$ independent measurements.
@@
whose uncertainty follows from the propagation of the uncertainties of $x_{i}$, i.e., $\sigma_{i}$, as
@@
For the ideal case of $x_{i}=\mu$ and $\sigma_{i}=\sigma$ (i.e., all measurements return the same value and the same uncertainty), we have $\mu^{\prime}=\mu$ and $\sigma_{\mu^{\prime}}=\sigma/\sqrt{N}$. Thus, as is well known, the rule of thumb is that with $N$ independent measurements of the same quantity the precision is improved by a factor of $\sqrt{N}$.
Given that the averaged PDF is not satisfactory, what would be the best way to extract the maximal information from the set of individual PDFs, or how to combine independent measurements of the same quantity? This is the question statisticians have asked, and have come up with reasonable ideas. One particular suggestion to follow here is the consolidation by conflation [^5] Throughout I will refer to this as “consolidation by conflation” or simply “consolidation”. due to T. P. Hill and J. Miller (Hill, 2011; Hill & Miller, 2011). In this approach the consolidated probability distribution of a variable $x$ is defined by
@@
Lacking general analytic results in general non-Gaussian situations and considering the unknown nature of the present problem, here an empirical approach is followed, i.e., the consolidated distribution (Equation (30)) is tried for mock individual PDFs and its property is experimentally investigated. Figure 4 shows examples of such a numerical experiment for the mock data sets of the two binaries shown in Figure 2. Remarkably, the consolidated distributions are close to Gaussian in both cases despite individual non-Gaussianities. As will be shown later, the (near) Gaussianity of the consolidated distribution is always satisfied empirically for any sufficiently large set of individual PDFs of mock or real binaries. Figure 4 indicates that the (near) Gaussianity is a consequence of the product (i.e., the conflation by Equation (30)) of the shapes of the individual PDFs (which were evenly sampled in $t/P$). As the right column of the figure shows this dramatic nature, although the majority of individual PDFs are biased toward the negative side (which is why the averaged PDF is also biased as discussed above), the minority rest is biased toward the positive side with narrower widths so as to make a “correct” balance with the majority.
In addition to being (nearly) Gaussian, the consolidated distribution correctly reproduces the input value of $\Gamma=0$. As indicated in Figure 4, the consolidated distributions give $\Gamma=0.005\pm 0.026$ and $-0.004\pm 0.042$ for the two sets of data. Note that the former value is more precise because the relative uncertainties of $v_{r}$ are smaller for the data of the binary with $a=0.8$ kau for the fixed uncertainty of $\sigma_{v_{r}}\approx 283$ m s-1. Although not shown here, precision can also be improved by increasing the sample size. These numerical experiments suggest that the consolidated distribution of $\Gamma$ from a sufficiently well-sampled individual PDFs have the potential to discriminate between Newtonian gravity and nonstandard gravities.
@@
In the present work, only the gravity anomaly parameter $\Gamma$ is of interest and can be consolidated for real data of wide binaries because the same gravity law is assumed in a certain acceleration range, but snapshot-observed wide binaries have different masses, orbital parameters, and inclinations, all of which are treated as nuisance parameters. However, the two specific sets of mock data shown in Figure 2 are special because each data set is composed of data for the same binary observed at sufficiently different times $t/P$, which of course cannot be realized in actual observations of wide binaries for obvious reasons. [^6] Not to mention, such a data set can be obtained for very close binaries with $P$ sufficiently small, e.g. several years. It is of some interest to consider the consolidation of the nuisance parameters for these mock data sets, not only as a matter of additionally testing the consolidation method, but also as a possible application of the method in very close binaries. Here I note that it is necessary to introduce and fit the rotation angle about the $z^{\prime}$-axis on the sky plane in general cases shown in Appendix A, if a PDF of the true inclination angle is desired. In the present case, it is not necessary because the rotation angle is zero by design in Figure 1. (In gravity tests with real binaries, the inclination angle is a nuisnace parameter and thus it is not necessary to introduce the rotation angle in the Bayesian modeling.)
@@
Similarly, Figure 6 shows that the individual PDFs of the eccentricity $e$ are not well shaped in general, but the consolidated distributions are well localized, nearly Gaussian, and reproduce the input value $e=0.75$ remarkably well.
Figure 7 shows the results for $\phi_{0}$. Because $\phi_{0}$ was allowed to vary freely in the range $-360^{\circ}<\phi_{0}<360^{\circ}$ and the input value was $-135^{\circ}$, we expect that individual PDFs will be localized around $-135^{\circ}$ or $225^{\circ}$. Most PDFs are indeed so, but there are a few exceptions. The most pronounced are the PDFs for $t/P=0.90$ and $0.95$, which correspond to the two points on the second quadrant of the $x^{\prime}y^{\prime}$ plane in Figure 2. Due to these outliers, the consolidated distributions (after transforming PDFs localized around $225^{\circ}$ by $-360^{\circ}$) are biased by $\approx-16^{\circ}$ and $\approx-22^{\circ}$, respectively for the two data sets of $a=0.8$ and $a=8$ kau, as can be seen in the lower panels. Compared with the net allowed range of $360^{\circ}$ the magnitude of these biases in $\phi_{0}$ is not large.
The bias in the PDFs of $\phi_{0}$ can be understood as follows. Equations (10) and (12) have dependence on $\phi_{0}$, but when their observational constraints given by Equations (9) and (6) are not strong enough, the prior constraint given by Equation (19) will play a major role. This means that some data near the periastron (that is, when $\phi\sim\phi_{0,\rm{true}}$) can lead to PDFs localized near $\phi_{0}\sim\phi\pm 180^{\circ}\sim\phi_{0,\rm{true}}\pm 180^{\circ}$ leading to bias. This appears to be the case for the two abovementioned outliers in each data set.
Thus, unlike $i$ and $e$, $\phi_{0}$ appears to be not perfectly recovered by the procedure for the above mock observations. A couple of relevant remarks need to be made here. First, as Figure 4 shows, the relatively small biases in $\phi_{0}$ do not affect the correct recovery of $\Gamma$ in both data sets. This can be understood by the fact that only Equation (10) has the direct dependence on $\Gamma$, and therefore the small bias in $\phi_{0}$ has little impact. Second, the bias is larger when the relative precision of the RVs is poorer, i.e., for the data set with $a=8$ kau. It is interesting to check whether the bias is reduced if the assumed uncertainties of the RVs are smaller for the same data. Figure 8 shows the PDFs of the four parameters for the mock binary with $a=8$ kau, but assuming 50 m s-1 for RV uncertainties so that $\sigma_{v_{r}}\approx\sqrt{2}\times 50\approx 71$ m s-1 for the relative RV $v_{r}$. The bias in $\phi_{0}$ is now reduced to $\approx-14^{\circ}$, while $\Gamma$ and the other parameters remain recovered correctly with smaller uncertainties as expected.
@@
Therefore, numerical experiments indicate that the procedure of Bayesian inferences with their consolidation can be used to measure $\Gamma$. Moreover, for a data set of one binary as in close binaries, the procedure can recover the parameters $e$ and $i$ correctly, and $\phi_{0}$ as well if the relative precision of velocities is good enough, which can be satisfied for close binaries. In close binaries, gravity is Newtonian (i.e., $\Gamma=0$ is fixed), so the procedure can be used to measure the total mass of the binary.
The above numerical experiments are designed to illustrate individual PDFs of parameters and develop/test the method of consolidation. For this reason, the mock data used in the above experiments have accurate values of velocities predicted by the Newtonian model. The assumed uncertainties of the velocities are only used to derive the PDFs of the model parameters. However, measured velocities will have errors due to measurement uncertainties, and thus errors need to be added to the model-predicted velocities for realistic simulations. For wide binary samples derived from the Gaia DR3 database to be considered in the next subsection, the projected components $v_{x^{\prime}}$ and $v_{y^{\prime}}$ have negligibly small errors, but the RV $v_{r}$ has large errors. Realistic simulations with the errors of $v_{r}$ added to the model-predicted values will be considered when the results of Bayesian modeling for real data are presented in Section 3.
@@ 2.3 Binary Sample
This study requires a sample of pure binaries that are free of hidden close components. Selecting and using pure binaries was originally suggested and considered by Hernandez et al. (2022) and Hernandez (2023) in the context of statistical tests of gravity with $v_{p}$. Here I apply a multistep screening process to ensure that no (or extremely few) kinematic contaminated systems survive. I start with the initial selection criteria that were used by Chae (2024a, b). The initial selection is based on the requirement that binaries in a bin of sufficiently high acceleration must agree with the expectation of Newtonian dynamics in a statistical sense, specifically in terms of the median or average properties. Such a selection (Chae, 2024a, b) can be achieved by requiring sufficiently good precision of PMs, parallaxes, and RVs for both components and dynamical consistency between them. Note that availability of Gaia-measured RVs (with a reasonable precision) particularly helps remove systems with hidden close companions because it implies a good single-star solution for the spectroscopic observations. However, as already acknowledged in Chae (2024a, b), there can be (a small number of) individual exceptions that elude such a statistical selection. [^7] Minor bumps in the high-velocity tails of normalized projected velocity distributions in Figures 15 - 17 of Chae (2024c) may indicate such exceptions.
Here I include some of $\delta_{A}<-28^{\circ}$ (hereafter $\delta$ refers to the decl. and the subscript $A$ refers to the brighter component of the pair) binaries as opposed to Chae (2023a, 2024a) who excluded the $\delta_{A}<-28^{\circ}$ data altogether (if distance is greater than 80 pc) because the dust extinction model adopted by Chae (2023a) did not cover that portion of the southern sky. I include those binaries with $\delta_{A}<-28^{\circ}$ that satisfy either distance $d_{A}<100$ pc or Galactic latitude $|b|>60^{\circ}$ as they are unlikely to be significantly affected by dust extinction (see Figure 3 of Chae 2023a). Another minor change is that I now select stars with $13.43.8<M_{G}<$ 13.4 (rather than $144<M_{G}<$ 14 where $M_{G}$ is the Gaia $G$-band absolute magnitude) in agreement with Figure 1 of Penoyre et al. (2022b).
@@
Note. (1) The base sample is defined as a sample with $f_{\rm{multi}}\rightarrow 0$ ($f_{\rm{multi}}$ is the fraction of hierarchical systems) in the Newtonian regime (it may include individual exceptions of hierarchical systems). This sample is revised and extended from Chae (2024a, b) to include some southern-sky binaries. (2) The base sample was selected from the El-Badry et al. (2021) catalog. Thus, it is free from flyby (chance alignment) and resolved triple (or higher-order) cases by their criteria. Here I use stricter criteria to remove any potentially suspicious systems as described in the text. Out of 4276 systems, there are 68 potential flybys and 35 triple+ systems. (3) The nominal sample is used to obtain nominal results, and the alternative sample and other samples with varied ruwe and $\sigma_{v_{r}}$ limits are considered to explore possible systematic uncertainties.
The above selection returns a sample of 4276 wide binaries, which is larger than the Chae (2024b) sample of 3557 wide binaries thanks to the inclusion of some of the southern-sky binaries. This sample will be referred to as the “base sample.” Table 1 gives a summary of the definition of the samples considered in this work, including the base sample. This sample is free of fly-bys (chance alignments) and resolved triples or higher-order multiples (to be referred to as triple+) by the criteria of El-Badry et al. (2021). Here I consider a stronger criterion to remove any potentially suspicious systems because individual modeling requires individually pure binaries. The neighboring space around each binary is defined by the following limits of the transverse and longitudinal separations:
@@
    the sky-projected displacement from the brighter component satisfies $\Delta R_{\rm{sky}}<\sqrt{20}s$ where $s$ is the projected separation between the pair, and
@@
I search the entire Gaia catalog of all objects within 200 pc downloaded from the Gaia DR3 data archive [^8] https://gea.esac.esa.int/archive/. For 35 systems out of the base sample, one or more stars are found beyond the known pair within the space. I identify these as potential resolved triple+ systems. For 68 systems, the fainter component of the pair as originally identified by El-Badry et al. (2021) is not found within the space defined above. I identify these as potential flybys. In total, 103 (2.4%) of the 4276 systems have potential kinematic contaminants. I flag these systems in the base sample and exclude them in the following sample selection.
@@
This work requires accurate and precise values for all three components of the relative velocity between the pair. This requirement is met for the two components of the projected velocity $v_{p}$ from the Gaia DR3 database: most projected velocities have uncertainties $\sigma_{v_{p}}\lesssim 35$ m s-1 as shown in Figure 9. Here $v_{p}$ and $\sigma_{v_{p}}$ are calculated using Equations (5) and (9) of Chae (2024a). RVs $v_{r}$ from the Gaia DR3 database have much larger uncertainties $\sigma_{v_{r}}>168$ m s-1. The uncertainty of $v=|\mathbf{v}|=\sqrt{v_{p}^{2}+v_{r}^{2}}$ ($\sigma_{v}$) is estimated from Monte Carlo distributions of $v_{p}$ and $v_{r}$ with their estimated uncertainties ($\sigma_{v_{p}}$ and $\sigma_{v_{r}}$) assuming normal distributions.
As $\sigma_{v_{r}}$ increases, $\sigma_{v}$ also increases and individual Bayesian modeling becomes increasingly less feasible and model parameters become more poorly constrained (and thus the results become less meaningful). As shown in Figure 2, wide binaries of low internal acceleration have relative velocities between the pair as low as a few hundreds meters per second or lower. Thus, for wide binaries of low internal acceleration, $\sigma_{v_{r}}$ significantly larger than a few hundreds meters per second are not likely to be useful. I put an initial hard cut $\sigma_{v_{r}}<500$ m s-1, which leaves only 652 binaries (15.6% of the base sample). This will be referred to as the “statistical sample” for Bayesian 3D modeling of this work.
Because reasonable precision is required for $v_{r}$ and a significant number of binaries is required for an unbiased inference of gravity (as will be shown below in Section 3), some compromise between the sample size and the tolerance of $\sigma_{v_{r}}$ is inevitable. Based on numerical experiments, I take $380$ m s-1 as a nominal choice for the upper limit of $\sigma_{v_{r}}$ but also consider $330$ m s-1 as an alternative choice. For $\sigma_{v_{r}}$ limits higher than $380$ m s-1, the bias caused by RV scatter seems to outweigh the increase in number statistics. For $\sigma_{v_{r}}$ limits lower than $330$ m s-1, the number statistics in the MOND regime is poor.
@@
The statistical sample with relatively precise $v_{r}$ values may still contain exceptional cases with kinematic contamination due to, e.g., unseen low-mass star(s) or massive Jovian planet(s) orbiting any component(s) of the binary system. A popular indicator of such close kinematic contamination is the Gaia renormalized unit weight error (ruwe) parameter [^9] https://gea.esac.esa.int/archive/documentation/. For general stars in the Gaia DR3 database a value significantly greater than $1.0$ (e.g., ruwe $>1.4$) could indicate that the source is nonsingle or otherwise problematic for the astrometric solution. Figure 10 shows the distribution of ruwe values for 1304 stars in the statistical sample; 92.2% (82.6%) satisfy ruwe $<1.4$ ($<1.2$). Nevertheless, it is possible that some of stars with ruwe $<1.4$ (or some other value) are unresolved nonsingle stars.
@@
Various simulations and observational studies suggest that ruwe values ranging from $1.15$ $1.6$ (e.g., Belokurov et al. 2020; Bryson et al. 2020; Fitton et al. 2022; Penoyre et al. 2022a; Castro-Ginard et. al. 2024) can help remove some (unresolved) close binaries. Since such a ruwe value does not provide a hard cut on close binaries (see, e.g., Figure 1 of Fitton et al. 2022) and the statistical sample is purer than the base sample (which is already pure in a statistical sense), I consider varying the upper limit of allowed ruwe values to explore the possible effects of the ruwe limit. For this purpose, I take $1.30$ as the nominal/standard choice for the upper limit of ruwe (see Figure 10). I consider other limits from ruwe $=1.1$ $2.0$ to explore possible systematic effects. Based on the nominal choices of both $\sigma_{v_{r}}$ and ruwe, I define the nominal/standard sample of 312 binaries as given in Table 1.